One type of jet engine has a 0.0001 probability of failure while in flight. For a jet that has four of these engines, what is the probability of at least two of them failing? Show all steps and explain how the answer was gotten
Answer #1

Let X denote the number of engines failing. Then, X~Binomial(n=4,p=0.0001)

P(X = r) = C) (0.0001)(1-0.0001)4-1,x = 0, 1, 2, 3, 4

Required probability =

P(X > 2) = P(X 2) + P(X = 3)+ P(X 4) = (. ) (0.0001)2(1 0.0001)4-2 + 5.9992 × 10-8 ) (0.0001 )3(1-0.0001)4-3 ) (0.0001 )4(1-0.0001 )4-4- 0.000000059992